# Collaborating with AI Agents: A Field Experiment on Teamwork, Productivity, and Performance

### v2

@@ Abstract
To uncover how AI agents change productivity, performance, and work processes, we introduce Pairit—an experimentation platform enabling humans and AI agents to collaborate in integrative workspaces. In a large-scale marketing experiment on the platform, 2310 participants were randomly assigned to human-human and human-AI teams. The teams exchanged 183,691 messages and created 63,656 image edits, 1,960,095 ad copy edits, and 10,375 AI-generated images while producing 11,138 ads for a large think tank. Analysis of fine-grained communication, collaboration, and workflow logs revealed that collaborating with AI agents increased communication by 63% and allowed humans to engage in 71% less direct text editing. While human-AI teams engaged in 18% more process and content communication, human-human teams engaged in 29% more social and emotional communication. Humans in human-AI teams experienced 73% greater productivity per worker and produced higher-quality ad copy, while human-human teams produced higher-quality images, suggesting AI agents require fine-tuning for multimodal workflows. Field tests of the ad campaigns accumulated $\sim 5\text{M}$ ad impressions and revealed that ads with higher image quality (produced by human-human collaborations) and higher text quality (produced by human-AI collaborations) performed significantly better on click-through rates, view through rates, and cost per click metrics. Together, these results suggest that human collaboration with AI agents significantly reshapes communication patterns and work processes and increases productivity, while improving some dimensions of output quality and deteriorating others. We hope the release of the extensible Pairit platform will accelerate RCTs of human-AI collaboration across a variety of work tasks and contexts.
@@ 1 Introduction
Artificial intelligence (AI) tools have garnered attention for their potential to improve productivity and performance (Eloundou et al., 2024, Bick et al., 2024). For example, large language models (LLMs) decreased the average time taken for mid-level professional writing tasks by 40% and increased output quality by 18% (Noy and Zhang, 2023). For job seekers, AI assistance with resumes increased job hiring by an average of 8% (Wiles et al., 2023), and for customer support workers, AI assistance increased productivity by an average of 14% (Brynjolfsson et al., 2023). Moreover, productivity gains were greater for lower-skilled workers (Noy and Zhang, 2023, Brynjolfsson et al., 2023, Choi and Schwarcz, 2023) and varied across task domains (Dell’Acqua et al., 2023). Evidence from an online labor market suggests there has already been a reduction in demand for freelance knowledge work with the advent of generative pre-trained transformer (GPT) models (Hui et al., 2023). In a study reviewing over 106 papers, Vaccaro et al. (2024) showed that human-AI groups outperformed humans alone in 85% of the studies. Our work aims to address two critical gaps in this burgeoning literature on AI’s impact on collaboration and productivity: Understanding 1) the role of multimodal AI agents (as distinct from LLM-enabled chatbots) and 2) the work process changes created by human-AI agent collaboration.
@@
While studies like Liu et al. (2023) explore levels of proactivity in AI agents, they typically do so without randomized controlled trials (RCTs) measuring productivity effects. The studies that use RCTs to estimate productivity effects tend to randomize access to LLM chatbots (e.g. Dell’Acqua et al. (2023) and Chen and Chan (2024)), which are not typically multimodal, do not include context, do not allow the chatbots to take independent actions or use APIs to call outside of the work environment, and do not provide a collaborative workspace where machines and humans can jointly manipulate output artifacts in real time. These innovations are meaningful because AI agents today possess all these features, yet the existing scientific literature studies none of them.
@@
Furthermore, we currently lack fine-grained task level insight into how human-AI collaboration changes work processes and communication patterns and how these changes affect productivity and performance. The vast majority of existing research focuses on the productivity effects of GPT chatbots on individual workers (Noy and Zhang, 2023, Dell’Acqua et al., 2023). Others investigate how AI changes people’s perceptions, beliefs, and behaviors (Tey et al., 2024, Costello et al., 2024). However, it is unclear how these interactions evolve in real-time collaborations, especially in environments where AI agents can take autonomous actions, adapt dynamically to human inputs, and participate in tasks requiring creativity and coordination. Current off-the-shelf experimental platforms and studies do not provide collaborative workspaces where researchers can precisely record and measure the collaboration itself: e.g., transcripts of messages between machines and humans, logs of edits to output artifacts, and API (application programming interface) calls to outside agents. In contrast, current AI applications, such as Notion AI, v0.dev, OpenAI Canvas, Cursor, and GitHub Copilot, already integrate AI agents into such collaborative workspaces and interfaces.
@@
To address these gaps, we developed Pairit, a novel experimental platform designed to study human-AI collaboration in real-world, extensible tasks. Pairit introduces several key innovations. It enables real-time collaboration between humans and AI agents, allowing participants to manipulate text, images, and workflows collaboratively in chat-enabled workspaces that mirror existing online AI collaboration work processes. The platform supports randomized pairings of humans and AI (i.e., human-human or human-AI teams) and allows for randomization of prompts and model fine-tuning. Critically, the AI can perform the complete set of equivalent actions that humans can perform in the collaborative workspace. These include sending chat messages, writing ad copy, editing ad copy, writing calls to action, editing calls to action, scrolling through images, editing images, selecting images, and generating new images using an external call to Dall-E 3. Moreover, Pairit captures every time-stamped keystroke, message, edit, swipe, scroll, selection, API call, and intermediate output, providing a rich dataset that allows for the detailed reconstruction of collaboration workflows. This represents a fundamental departure from the existing literature, which enables RCT-based evaluation of the outputs and productivity implications of human collaboration with LLM-based chatbots and co-pilots, but does not enable randomized experiments analyzing the task-level productivity and work process changes created in human collaborations with fully functioning, multimodal AI agents.
@@
To test the Pairit platform and to study the work process, productivity, and performance implications of human collaboration with fully functional AI agents, we conducted a large-scale randomized study of the effects of human-AI collaboration on advertising design and creation—a task requiring creativity, iteration, and precision. A total of 2,310 participants, representative of the U.S. population, were recruited through Prolific and randomly paired into human-human or human-AI teams. Teams worked collaboratively to create marketing campaigns for a think tank’s year-end annual report, including generating and selecting ad images and writing ad copy and calls to action. This process was fully recorded, resulting in a dataset including 11,138 ads, 183,691 messages, 1,960,095 text edits, 63,656 image edits, and 10,375 AI-generated images, offering an unprecedented level of detail with which to understand work processes, teamwork dynamics, communication, productivity and output quality.
@@
Once the lab portion of the experiment was completed, we conducted a field experiment on the ads produced by human-human and human-AI teams. We obtained human and AI quality ratings of the ads, including the quality of the ad copy and images, and the human- and AI-evaluated likelihood of consumer engagement with the ads (measured by click-through rates). We then ran the ads a real online display ad ecosystem, generating over 4.9 million impressions, and evaluated click-through rate, cost-per-click, view-through rate and view through duration metrics on the annual report using the platform’s ads API and DocSend’s view metrics, which allow us to record how much of the report consumers read, page by page, after clicking through on the ads.
In mining the rich collaboration data to understand how AI changes team communication, we found individuals in human-AI teams sent 63% more messages than those in human-human teams and that results were consistent at the team level. Furthermore, both the humans and the AI in human-AI teams sent 18% more content- and process-oriented messages than those in human-human teams, especially messages containing suggestions, instructions, prioritization, and planning. Conversely, humans in human-human teams sent 29% more social and emotional messages, including more messages that expressed rapport building, self-assessment, and concern, than humans in human-AI teams. These results demonstrate that human-human teams spent considerable time and effort on social and emotional calibration and regulation, while human-AI teams were task- and process-oriented.
We also tracked work process changes in the collaborative workspace where humans and agents could both edit the ad copy as well as generate images using external APIs (i.e., Dall-E 3). We found that human-AI teams made 71% fewer edits to the copy due to AI’s proficiency in writing high-quality ad copy. In contrast, the AI was worse at predicting image quality, thus disadvantaging human-AI teams in the visual dimensions of ad creation and leading to human-AI teams creating lower-rated images. These differences in collaboration outcomes indicate a shift in workload that allowed human participants in human-AI teams to engage more in content generation with less social coordination, like rapport building, than humans in human-human teams.
These communication, collaboration, and work process changes then translated into measurable differences in productivity and performance. Humans in human-AI teams experienced 73% greater productivity per worker and human evaluations scored the ads created by human-AI teams higher on text quality but lower on image quality. AI evaluations scored the ads created by human-AI teams higher on text quality and equally on image quality. Field tests of the ad campaigns revealed that ads with higher image quality (produced more often by human collaborations) and higher text quality (produced more often by human-AI collaborations) performed significantly better on click-through rates, view through duration and cost per click metrics. As human-AI teams produced higher quality text and lower quality images, human-AI ads performed similarly to those created by human-human teams overall. Together, these results suggest that human collaboration with AI agents significantly reshapes communication patterns and work processes and increases productivity, while improving some dimensions of output quality and deteriorating others.
@@
Although prior studies have shown that AI tools can improve productivity and reduce task completion times (Noy and Zhang, 2023, Brynjolfsson et al., 2023), AI is often treated as a passive tool rather than as an active collaborator. As AI agents become integral to modern workflows, researchers are beginning to explore their role as work collaborators, rather than mere tools, emphasizing the importance of trust, transparency, and integration in human-AI partnerships (Makarius et al., 2020, Anthony et al., 2023, Collins et al., 2024). Our work contributes to this emerging literature by presenting the first task-level randomized experiment measuring the work process, productivity and performance implications of human-AI collaboration with fully functional, multimodal AI agents. We hope the release of the extensible Pairit platform will accelerate the deployment of RCTs of human-AI collaboration across a variety of work tasks and contexts in future research.

### v3

@@ Abstract
We examined the mechanisms underlying productivity and performance gains from AI agents using a large-scale experiment on Pairit, a platform we developed to study human-AI collaboration. We randomly assigned 2,234 participants to human-human and human-AI teams that produced 11,024 ads for a think tank. We evaluated the ads using independent human ratings and a field experiment on X which garnered $\sim$5M impressions. We found human-AI teams produced 50% more ads per worker and higher text quality, while human-human teams produced higher image quality, suggesting a jagged frontier of AI agent capability. Human-AI teams also produced more homogeneous outputs (or diversity collapse) as their ads were more self-similar. The field experiment revealed higher text quality (produced by human-AI teams) improved click-through rates and view-through duration, while higher image quality (produced by human-human teams) improved cost-per-click rates. We found three mechanisms explained these effects. First, human-AI collaboration was significantly more task-oriented, with 25% more task-oriented messages and 18% fewer interpersonal messages. Second, human-AI collaboration displayed more delegation, as participants delegated 17% more work to AI agents than to human partners and performed 62% fewer direct text edits when working with AI. Third, recognition that the collaborator was an AI moderated these effects as participants who correctly identified they were working with AI were more task-oriented and more likely to delegate work. These mechanisms, in turn, explained performance as task-oriented communication improved ad quality, specifically when working with AI, while interpersonal communication reduced ad quality; delegation improved text quality but had no effect on image quality (consistent with the jagged frontier) and was positively associated with diversity collapse, creating homogeneous outputs of higher average quality. The results suggest AI agents drive changes in productivity, performance, and output diversity by reshaping teamwork.
@@ 1 Introduction
Artificial intelligence (AI) tools have garnered attention for their potential to improve productivity and performance (Eloundou et al., 2024, Bick et al., 2024). For example, large language models (LLMs) decreased the average time taken for mid-level professional writing tasks by 40% and increased output quality by 18% (Noy and Zhang, 2023). For job seekers, AI assistance with resumes increased job hiring by an average of 8% (Wiles et al., 2023), and for customer support workers, AI assistance increased productivity by an average of 14% (Brynjolfsson et al., 2023). Moreover, productivity gains were greater for lower-skilled workers (Noy and Zhang, 2023, Brynjolfsson et al., 2023, Choi and Schwarcz, 2023) and varied across task domains (Dell’Acqua et al., 2023). Evidence from an online labor market suggests there has already been a reduction in demand for freelance knowledge work with the advent of generative pre-trained transformer (GPT) models (Hui et al., 2023). In a study reviewing over 106 papers, Vaccaro et al. (2024) showed that human-AI groups outperformed humans alone in 85% of the studies.
@@
While studies like Liu et al. (2023) explore levels of proactivity in AI agents, they typically do so without randomized controlled trials (RCTs) measuring productivity effects. The studies that use RCTs to estimate productivity effects tend to randomize access to LLM chatbots [e.g. Dell’Acqua et al. (2023) and Chen and Chan (2024)], which are not typically multimodal, do not include context, do not allow the chatbots to take independent actions or use APIs to call outside of the work environment, and do not provide a collaborative workspace where machines and humans can jointly manipulate output artifacts in real time. These innovations are meaningful because AI agents today possess all these features, yet the existing scientific literature studies none of them.
@@
Furthermore, we currently lack fine-grained task level insight into how human-AI collaboration changes work processes and communication patterns and how these changes affect productivity and performance. The vast majority of existing research focuses on the productivity effects of GPT chatbots on individual workers (Noy and Zhang, 2023, Dell’Acqua et al., 2023) or how AI changes people’s perceptions, beliefs, and behaviors (Tey et al., 2024, Costello et al., 2024). However, it is unclear how these interactions evolve in real-time collaborations, especially in environments where AI agents can take autonomous actions, adapt dynamically to human input, and participate in tasks requiring creativity and coordination. Our lack of evidence on in vivo human-AI collaboration exists, in part, because off-the-shelf experimental platforms and studies do not provide collaborative workspaces where researchers can precisely record and measure the collaboration itself: e.g., transcripts of messages between machines and humans, logs of edits to output artifacts, and API (application programming interface) calls to outside agents or tools. In contrast, current AI applications, such as Notion AI and Cursor, already integrate AI agents into such collaborative workspaces and interfaces.
@@
To address these gaps, we developed Pairit, a novel experimentation platform designed to study human-AI collaboration in real-world, extensible tasks. Pairit introduces several key innovations. It enables real-time collaboration between humans and AI agents, allowing participants to manipulate text, images, and workflows collaboratively in chat-enabled workspaces that mirror existing online AI collaboration work processes. The platform supports randomized pairings of humans and AI (i.e., human-human or human-AI teams) and allows for randomization of prompts and model fine-tuning. Critically, the AI can perform the complete set of equivalent actions that humans can perform in the collaborative workspace. In the marketing experiment analyzed in this study, these include sending chat messages, writing ad copy, editing ad copy, writing calls to action, editing calls to action, scrolling through images, editing images, selecting images, and generating new images using an external call to Dall-E 3. Moreover, Pairit captures every time-stamped keystroke, message, edit, swipe, scroll, selection, API call, and intermediate output, providing a rich datasets that allow for the detailed reconstruction of collaboration workflows. This represents a fundamental departure from the existing literature, which enables RCT-based evaluation of the outputs and productivity implications of human collaboration with LLM-based chatbots and co-pilots, but does not enable randomized experiments analyzing the task-level productivity and work process changes created in human collaborations with fully functioning, multimodal AI agents.
@@
To study the work process, productivity, and performance implications of human collaboration with multi-modal AI agents, we conducted a large-scale randomized study of human-AI collaboration on advertising design and creation, a task requiring creativity, iteration, and precision. A total of 2,234 participants, representative of the U.S. population, were recruited through Prolific and randomly assigned to human-human or human-AI teams using the Pairit platform. Teams worked collaboratively to create marketing campaigns for a think tank’s year-end annual report, including generating and selecting ad images and writing ad copy and calls to action. This process was fully recorded, resulting in a dataset including 11,024 ads, 182,607 messages, 1,889,559 text edits, 62,119 image edits, and 10,074 AI-generated images, offering an unprecedented level of detail with which to understand work processes, communication, productivity and output quality.
@@
Once the lab portion of the experiment was completed, we conducted a field experiment on the ads produced by human-human and human-AI teams. We obtained human and AI quality ratings of the ads, including the quality of the ad copy and images, and the human- and AI-evaluated likelihood of consumer engagement with the ads (measured by click-through rates). We then ran the ads in a real online display ad ecosystem, generating over 4.9 million impressions on X, and evaluated click-through rate, cost-per-click, view-through rate, and view-through duration metrics on the annual report using the platform’s ads API and DocSend’s view metrics, which allow us to record how much of the report consumers read, page by page, after clicking through on the ads.
Our analyses examine three performance effects documented in the AI literature but in an agentic collaboration context: productivity and performance (Noy and Zhang, 2023, Brynjolfsson et al., 2023), the jagged frontier of AI capabilities across tasks (Dell’Acqua et al., 2023, Gans, 2026), and diversity collapse, or the homogenization of AI-assisted outputs (Padmakumar and He, 2024, Chen and Chan, 2024, Hao et al., 2026). We found evidence of all three effects in our agentic setting: human-AI teams produced 50% more ads per worker, higher-quality text, and more homogeneous outputs, while human-human teams produced higher-quality images. Field tests revealed that higher image quality (from human-human teams) improved cost-per-click while higher text quality (from human-AI teams) improved click-through rates and view-through rates, with these offsetting quality effects explaining similar overall field experiment ad performance across team types.
Most importantly, we found that specific teamwork dynamics—task-orientation, delegation, and AI recognition—moderate these effects. While our treatment randomizes AI agents at the extensive margin (humans working with or without AI), the data we collected on communication patterns, delegation decisions, and AI recognition capture how people interact with, collaborate with, and delegate work to AI agents. In mining the rich collaboration data produced by Pairit, we found individuals in human-AI teams sent 62% more messages than those in human-human teams. Furthermore, human-AI teams sent 25% more content- and process-oriented messages, especially messages containing suggestions, instructions, prioritization, and planning. Conversely, human-human teams sent 18% more social and emotional messages, including messages that expressed rapport building, self-assessment, and concern. Task-orientation, when working with AI, was then associated with higher text quality, image quality and likelihood of clicking, while interpersonal communication was associated with lower text quality, image quality and likelihood of clicking. We also tracked work process changes: human-AI teams made 62% fewer direct edits to the copy and delegated 17% more work to their AI partners. Higher delegation was then associated with improved text quality but not image quality, and with reduced output diversity. To support the idea that AI reshapes teamwork, we found that participants who correctly recognized that they were working with AI were more task-oriented and more likely to delegate work to their AI partners, teamwork processes associated with higher performance. Together, these results strongly suggest that collaboration with AI agents reshapes work processes in specific ways and that these new teamwork dynamics moderate productivity gains and jagged frontier quality effects in human-AI collaboration.
@@
Although prior studies have shown that AI tools can improve productivity and reduce task completion times (Noy and Zhang, 2023, Brynjolfsson et al., 2023), AI is often treated as a passive tool rather than as an active collaborator in prior research. As AI agents become integral to modern workflows, researchers are beginning to explore their role as work collaborators, rather than mere tools, emphasizing the importance of trust, transparency, and integration in human-AI partnerships (Makarius et al., 2020, Anthony et al., 2023, Collins et al., 2024). Our work contributes to this emerging literature by presenting the first task-level randomized experiment measuring the work process, productivity and performance implications of human-AI collaboration with fully functional, multimodal AI agents. The Pairit platform serves as the methodological infrastructure enabling this comparison, and the field experiment provides external validation of quality and performance differences. The core contribution is the empirical analysis of how human-AI collaboration reshapes task-orientation, delegation, communication, and recognition, as well as output quality, performance, and productivity. We hope these observations inform future research as AI agents become integral to the workplace.
2 Theory
2.1 Teamwork in Human-Human and Human-AI Teams
Effective teamwork requires more than individual competence and depends critically on how team members coordinate, communicate, and manage interpersonal relationships (Marks et al., 2001). These social dynamics generate collective intelligence that exceeds the sum on individual abilities (Woolley et al., 2010). A foundational distinction in this literature separates taskwork, the technical work itself, from teamwork, the processes that enable individuals to work together effectively (Salas et al., 1992). Teamwork encompasses three types of processes: transition processes where teams plan and set goals, action processes where they coordinate and monitor progress, and interpersonal processes where they manage conflict and build motivation (Marks et al., 2001). Teams cycle through these phases as they execute work.
Human teams invest substantial effort in all three. They build shared mental models through ongoing communication about goals, strategies, and progress, enabling them to anticipate each other’s actions (Mathieu et al., 2000). This shared understanding allows teams to negotiate who does what and when. Members also build trust, provide support, and resolve conflict to maintain cohesion, though this consumes time and cognitive resources.
The emergence of AI agents that can act autonomously within collaborative workspaces represents a fundamental shift from AI as a tool to AI as a teammate. Unlike chatbot-style interfaces where humans initiate each interaction, agents in our setting have full observability of the shared workspace and can read, write, and edit artifacts just as a human teammate would (Wang et al., 2024). This shift raises fundamental questions about how teamwork dynamics change when one team member is artificial (Seeber et al., 2020). Human-robot interaction offers relevant insights, though most of this work examines physical tasks with clearly delineated roles. Even so, consistent patterns emerge: in manufacturing, healthcare, and military contexts, people adapt their behavior based on perceived robot capabilities and reliability (Hancock et al., 2011), and trust calibration proves central to effective collaboration (Lee and See, 2004, Parasuraman and Manzey, 2010). Poorly calibrated trust leads to either over-reliance on flawed automation or underutilization of capable systems. These dynamics can enable human-AI combinations to outperform either alone, though the magnitude depends on task characteristics and collaboration design (Vaccaro et al., 2024, Collins et al., 2024). Knowledge work introduces additional complexity: AI agents can perform the same cognitive tasks as human teammates, creating ambiguity about optimal task allocation. At the organizational level, AI integration affects not just dyadic collaboration but broader structures of coordination and role design (Bankins et al., 2024).
The shift from human to AI teammates may alter the fundamental nature of collaboration. For example, human teammates bring idiosyncratic knowledge, experiences, and perspectives that create cognitive diversity (Hong and Page, 2004). This diversity allows teams to outperform homogeneous groups of higher-ability individuals, but only when social dynamics like turn-taking and mutual responsiveness allow diverse perspectives to surface (Woolley et al., 2010). These same social dynamics carry expectations of reciprocity, fairness, and recognition. AI teammates, by contrast, bring consistent availability and (within their capabilities) reliable execution, but without the diversity that emerges from different life experiences or the social fabric that binds human teams. We examine how these differences manifest in three process mechanisms: communication patterns, delegation behavior, and recognition of partner identity.
2.2 Task-Oriented vs Interpersonal Communication in Human-AI Teams
Human-human collaboration requires social maintenance, including building rapport, providing emotional support, and repairing conflicts when they arise (Marks et al., 2001, Hoegl and Gemuenden, 2001). These interpersonal processes are not merely overhead; they build the trust and psychological safety that enable effective coordination (Edmondson, 1999). Psychological safety, a shared belief that the team is safe for interpersonal risk taking, reduces the vulnerability that prevents team members from admitting errors or proposing untested ideas. Yet these processes also consume time and cognitive resources that could otherwise be devoted to task execution.
AI partners require none of this social maintenance. They do not take offense, need encouragement, or require relationship repair after disagreements. This absence of social needs may fundamentally shift communication patterns and change the nature of optimal communication for productivity and performance in human-AI teams. When AI agents can take actions on shared work, humans may shift from direct task execution to directing work through communication, adopting a more instrumental stance than they would with human partners (Tey et al., 2024). We therefore examine whether communication with AI teammates becomes more task-oriented and less interpersonal.
Communication patterns may also affect performance outcomes, not just differ across conditions. Team communication predicts performance, though the strength of this relationship depends on task characteristics and team composition (LePine et al., 2008). One reason is that interpersonal communication builds trust and enables implicit coordination, where members anticipate each other’s needs without explicit discussion (Rico et al., 2008). Yet these benefits come at a cost. Attention is a limited resource (Kahneman, 1973), and social maintenance consumes cognitive capacity that could otherwise support task execution. For tasks with clear objectives where implicit coordination matters less than explicit direction, redirecting attention from relationship management to task completion may improve productivity.
Because AI teammates do not require negotiation or consensus-building, task-oriented communication may prove more effective with AI than with humans. Conversely, interpersonal communication that builds rapport with human teammates may be counterproductive with AI partners, consuming attention without yielding the trust and coordination benefits it provides in human teams. We examine here whether task-oriented communication improves performance while interpersonal communication harms performance in human-AI teams.
2.3 Delegation in Human-AI Teams
When working with capable partners, individuals face choices about task allocation, deciding what to do themselves versus what to delegate. In human teams, social considerations shape these choices. Norms of fairness discourage assigning disproportionate workloads, concerns about overburdening partners limit requests, and expectations of reciprocity create implicit debts. Economic theory identifies a related constraint. Team production creates moral hazard because individual contributions are difficult to observe (Holmström, 1982). Delegation in human teams thus involves interpersonal negotiation and trust, not just task optimization.
AI teammates may alter these dynamics. Without relationships and social standing to maintain and protect, delegation to AI carries fewer interpersonal costs. Prior work shows people adjust their behavior with AI. For example, people respond differently to AI and human recommendations, exhibiting algorithm aversion in some contexts and appreciation in others (Dietvorst et al., 2015, Logg et al., 2019), and they reduce effort when AI predictions are confident (Agarwal et al., 2025, Fügener et al., 2022). Yet most of this work examines AI as advisor or predictor rather than as a collaborator taking actions on shared work. We examine delegation in settings where AI can take actions directly, editing text, selecting images, and communicating with human partners. Without the social costs of delegating to a peer, humans may delegate more readily to an AI agent. This shift from doing to delegating represents a fundamental change in how work is accomplished, from direct execution to supervision and direction (Morgeson et al., 2010, Seeber et al., 2020). We therefore examine whether humans delegate more work to AI collaborators than to human collaborators.
Beyond shifting how work is accomplished, delegation may affect output quality. Delegation effectiveness depends on the delegate’s competence (Leana, 1986). This principle extends to AI, which exhibits a jagged frontier of capabilities, excelling at some tasks while struggling with others (Dell’Acqua et al., 2023). We examine delegation effects across this frontier. Large language models are trained extensively on text and instruction-following, suggesting delegation may benefit text quality. Visual selection requires predicting audience response, a task where AI capabilities are less established, so delegation effects on image quality may be weaker or absent.
Delegation also has downstream consequences for output characteristics. More delegation means more of the final output reflects the partner’s contributions. When that partner is AI, delegation amplifies both the benefits of AI collaboration (Noy and Zhang, 2023, Brynjolfsson et al., 2023) and its costs, including homogeneity (Padmakumar and He, 2024) and limitations outside AI capabilities (Dell’Acqua et al., 2023). We therefore examine whether delegation is associated with improved output quality and whether these benefits vary with AI capability.
2.4 AI Partner Recognition
The task-oriented communication and delegation patterns that we argue improve human-AI performance assume that participants recognize their partner as AI. Yet people automatically apply social rules to computers (Nass and Moon, 2000, Reeves and Nass, 1996), defaulting to politeness, turn-taking, and concerns about fairness even when interacting with machines. This tendency may be counterproductive when it leads to social investment that yields no benefit with AI partners. Recognition may enable collaborators to override these defaults and match their behavior to the actual demands of the interaction.
Perceived partner identity affects trust, communication, and delegation patterns (Zhang et al., 2023, Glikson and Woolley, 2020), and appropriate reliance on automation depends on accurate assessment of a partner’s capabilities (Lee and See, 2004). Human-appropriate collaboration strategies include building consensus before acting, softening feedback to preserve relationships, and investing in rapport that enables future reciprocity. These strategies make sense with human partners who have feelings, memories, and participate in ongoing relationships, but they also consume effort without benefit when applied to AI. Participants who recognize their AI partner can instead adopt task-oriented communication (Section 2.2) and delegate more readily (Section 2.3), iterating quickly without social cost and focusing on task outcomes rather than relationship maintenance. Those who mistakenly believe their AI partner is human may achieve suboptimal outcomes not because AI collaboration is inherently limited, but because they apply strategies suited to a human partner rather than an AI partner.
Recognition may therefore improve performance by enabling behavioral adaptation. In our design, participants are not told whether their partner is human or AI until after the task, allowing us to measure the effects of the actual and perceived identity of the partner. We therefore examine whether recognition of an AI partner predicts adoption of the collaboration behaviors that we argue enhance human-AI performance, and whether these behaviors, in turn, affect output quality.